# Quantifying the generation of negatively charged boron vacancies in He-ion irradiated hexagonal boron nitride

### v1

@@ III.3 Molecular dynamics simulation of vacancy creation
In a next step, we correlate the experimentally found production yield to molecular dynamics (MD) simulations of the ion matter interaction. We model the 30 keV He-ions impinging onto a -thick hBN sheet. The calculations are performed for a reference fluence (see Appendix A.6) and incidence angle. Contrary to the binary collision approximation (BCA), used in common SRIM (stopping range of ions in matter) simulations [21, 41, 19, 28, 17], MD simulations take into account the molecular structure of the target materials (an example is shown in Figure 4 c), which yields more precise estimates of the defect production. Figure 4 a depicts the depth profile of nitrogen (blue) and boron (red) vacancies at the reference fluence corresponding to . On average, we find densities of and for nitrogen and boron, respectively. Figure 4 b compares the MD approach to simulations based on BCA. For the latter we find on average higher defect densities, which agrees with the general notion that these simplified methods tend to overestimate the defect creation.
@@
We thus compare the results from the MD simulations and from the microscopic charge model to estimate the production efficiency of the irradiation process. The results are shown in the last column of Table 1. Overall, our analysis shows that the production efficiency is rather low at about 0.2 %, implying that although with larger implantation fluences more vacancies are created, a large fraction of vacancies does not get promoted to V–B and remains optically inactive [67, 68, 44]. It should be pointed out that our simulations do not account for in situ annealing of defects through annihilation of vacancies with interstitials, which are mobile at room temperature [69, 70],as it happens at a macroscopically long time scale. The annealing should decrease the number of vacancies, independently of their charge state. Our values for the defect yield constitute a lower bound, since we included the potential effect of a constant background charge. In contrast, Ref. 44 reported a larger negatively charged V–B fraction of % for a fluence , which is comparable to our lowest irradiation fluence for which ODMR was measured. However, we note that the latter study concludes that the defect yield decreases steeply with ion fluence. With our background correction, the derived defect yield is constant within the experimental uncertainty. In other words, we tentatively state that the microscopic charge model without background correction tends to significantly overestimate the V–B density at low ion fluences.
@@ IV Conclusions
In summary, we quantified the defect creation within homogeneous hBN films using a helium ion microscope to site-selectively pattern of V–B defect ensembles. We analyzed the optical properties of the emitters as a function of ion fluence over more than two orders of magnitude. The spatially precise generation allowed us to study the magnetic response probed by ODMR contrast measurements in a consistent and systematic way. We simulated the level splitting obtained experimentally from the ODMR spectra employing a microscopic charge model, where the spatially random distribution of local charges results in a level splitting in ODMR. As a main finding, we observe a discrepancy between the scaling of the charge density, or equivalently the level splitting in ODMR, and the scaling of the defect PL, or equivalently the density of optically active defects, as function of ion fluence. Unlike other studies, which attributed the charge density in the microscopic model solely to V–B defects, we propose to include a background density in the model. With this modification, we were able to estimate the defect creation yield and compare it to molecular dynamics simulations. Our results systematically quantify the He-ion based creation of V–B defects in hBN. On the one hand, further developments may involve implementing new theoretical models with the aim of addressing the fundamental question of the origin of the V–B charge [71], such as the nature of the charged state (dynamical or static) and the role of the local defect environment, as has been done for N centers [72]. On the other hand, spatially precise creation of such emitters as demonstrated here may enable their integration in gated, light-emitting hetero-stacks, in which a tunneling current can be coupled to the in-gap defect states by means of graphene electrodes [56, 73].
@@ Acknowledgements.
N. S. thanks the Novo Nordisk Foundation NERD Programme (project QuDec NNF23OC0082957). Work at the Technical University of Munich was supported through the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Munich Center for Quantum Science and Technology No. (MCQST)-EXC-2111-390814868. C.K., L.G., J.F. and A.W.H. acknowledge funding through the International Graduate School of Science and Engineering (TUM-IGSSE, project BrightQuanDTUM). N.S. and M.W. acknowledge the support from the Independent Research Fund Denmark, Natural Sciences (project no. 0135-00403B) and from the Danish National Research Foundation through NanoPhoton - Center for Nanophotonics (project no. DNRF147). A.H. and I.B. acknowledge the NNF Biomag project (NNF21OC0066526). A.V.K. acknowledges funding from the German Research Foundation (DFG), projects KR 4866/9-1, and the collaborative research center “Chemistry of Synthetic 2D Materials” CRC-1415-417590517. Generous CPU time grants from the Technical University of Dresden computing cluster (TAURUS) and Gauss Centre for Supercomputing e.V., Supercomputer HAWK at Höchstleistungsrechenzentrum Stuttgart are greatly appreciated.Conflict of Interest
@@
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
@@ A.2 Optical Characterization
Optical properties are characterized by diffraction-limited confocal scanning photoluminescence (PL) mapping, Raman microscopy, and time-resolved photoluminescence (TRPL) under ambient conditions. PL measurements are performed on a custom confocal microscope featuring a 532-nm laser (Cobolt, Hübner GmbH). The objective is a Nikon TU Plan Fluor /0.9 NA EPI D. The signal is collected by a spectrometer (Andor-Solis SR-303i, 150 grooves/mm grating with blaze at 800 nm) equipped with a CCD (Andor Newton). Before the He-ion beam irradiation described below, we verify that the selected hBN flakes contain low background luminescence in the spectral range where emission from V–B is expected ( – ) as well as that only a few intrinsic quantum emitters, with typical emission lines in the visible, are present. Raman spectroscopy is carried out in a commercial Raman microscope (WiTec Alpha, 532 nm laser, cw-excitation, 300 grooves/mm grating with blaze at 500 nm, Zeiss EC Epiplan-Neofluar objective lens / 0.9 NA).
Time-resolved PL measurements (Figure A.3) are performed on the same setup as for PL experiments. For the excitation, we use a PicoQuant 515 nm laser diode in pulsed mode (200 µW of power, 10 MHz repetition rate), in combination with an avalanche photodiode (MPD) and a Time Tagger 20 (Swabian Instruments). The width of the instrument response function (measured on a reflective substrate) is around 250 ps. For fitting, we use the Python package Lifefit [74].
@@ A.3 Helium-ion irradiation
To generate luminescent defects in the hBN films, we use a focused He-ion beam in a helium-ion microscope (Zeiss Orion). As irradiation patterns, we choose arrays of circular patches with diameter and 2 µm pitch. The former is on the order of the typical spot size in the confocal microscope, such that He-irradiated spots appear point-like in the PL images (Figure A.2), and the latter is much larger, so that the exposure spot can be well separated laterally. We use an acceleration voltage of , a beam spacing of 5 nm in x- and y-directions, a beam aperture of , dwell time of and a beam current of . The values of the irradiation fluence are set in on the HIM (see Table 1 in the main text for the fluences employed). To minimize unwanted irradiations, we use a large field of view and low imaging fluence for locating the heterostructures. For hBN placed directly on Si/SiO2, we find significant background PL after ion exposure. The PL originates from both the Si/SiO2 substrate and the hBN (Figure A.4). To avoid such background PL, all the measurements in the main manuscript are conducted on hBN/graphite/SiO2/Si heterostructures, resulting in spectra virtually free from unwanted background luminescence.
@@ A.6 Molecular Dynamics (MD) simulations
MD simulations are carried out using the LAMMPS package [58] to extract the statistics of defect creation produced by the interaction of He ions with hBN. The interaction of the ion beam with the hBN structure is modeled via an extended Tersoff potential for boron nitride (BN-ExTeP) [76], with a smooth transition to the Ziegler–Biersack–Littmark (ZBL) potential [77] for small distances. The simulation spans a material thickness of and is performed for a reference fluence, given by a number of 990 ions and a covered surface of 35 µm 35 µm. The MD approach is compared to other Monte Carlo methods, which are the TRIDYN code [78] and the SRIM software [79].

### v3

@@ III.3 Molecular dynamics simulation of vacancy creation
In a next step, we correlate the experimentally found production yield to molecular dynamics (MD) simulations of the ion matter interaction. We model the 30 keV He-ions impinging onto a -thick hBN sheet. The calculations are performed for a reference fluence (see Appendix A.6) and incidence angle. Contrary to the binary collision approximation (BCA), used in common SRIM (stopping range of ions in matter) simulations [21, 41, 19, 28, 17], MD simulations take into account the molecular structure of the target materials (an example is shown in Figure 4 c), which yields more precise estimates of the defect production. Figure 4 a depicts the depth profile of nitrogen (blue) and boron (red) vacancies at the reference fluence corresponding to . On average, we find densities of and for nitrogen and boron, respectively. Figure 4 b compares the MD approach to simulations based on BCA. For the latter, we find on average higher defect densities, which agrees with the general notion that these simplified methods tend to overestimate the defect creation.
@@
We thus compare the results from the MD simulations and from the microscopic charge model to estimate the production efficiency of the irradiation process. The results are shown in the last column of Table 1. Overall, our analysis suggests that the production efficiency is rather low at about 0.2 %, implying that, although with larger implantation fluences more vacancies are created, a large fraction of vacancies does not get promoted to V–B and remains optically inactive [71, 72, 44]. However, the actual fraction of defects that become optically active may be substantially better, as our simulations do not account for in situ annealing of defects through annihilation of vacancies with interstitials, which are mobile at room temperature [73, 74]. The annealing should decrease the number of created vacancies, independent of their charge state. Unfortunately, direct simulation of annealing process using MD is challenging, as annealing happens at macroscopically long time scale, unachievable using MD simulations. In addition, charge transfer (e.g., between vacancies and interstitials) and mutual attraction of the defects through screened Coulomb interaction are possible, which should further affect the defect dynamics. The only way to model this process is to first perform extensive DFT calculations to assess all migration barriers, accounting for the interaction between the defects, and then perform kinetic Monte Carlo simulations. Such simulations will require substantial computational effort and are beyond the scope of this work. Overall, our values for the defect yield constitute a lower bound, since we cannot quantitatively account for the impact of defect annealing and since we included the potential effect of a constant background charge in the analysis of the ODMR splitting. In contrast, Ref. 44 reported a larger negatively charged V–B fraction of % for a fluence , which is comparable to our lowest irradiation fluence for which ODMR was measured. However, we note that the latter study concludes that the defect yield decreases steeply with ion fluence. With our background correction, the derived defect yield is constant within the experimental uncertainty. In other words, we tentatively state that the microscopic charge model without background correction tends to significantly overestimate the V–B density at low ion fluences.
In this context, one may ask how the quality of the hBN crystal will impact the measured OMDR splitting and therefore the corresponding, extracted background charge. For this purpose, we compare our ODMR results with previously reported studies that explicitly observed V–B defects and provided sufficient information about the hBN quality, the irradiation parameters, and the ODMR results (Table 3). Overall, different studies, using different irradiation methods (both ions and neutrons) and different source crystals, report comparable values of the V–B ODMR splitting (about 100 - 150 MHz). Furthermore, a consistent observation throughout is that lower ion fluences result in smaller splitting, in line with our results. The lowest values (about 90 MHz) were achieved in high-pressure, high-temperature grown hBN (from NIMS, Japan) [50, 52]. Indeed, such high-pressure, high-temperature grown hBN is generally considered to be less defective, as evidenced for example by its lower intrinsic defect-related luminescence [75], compared to hBN crystals provided by commercial sources, such as HQgraphene (as used in the present work). Yet, we point out that based on the currently available data, we cannot make a conclusively statement on the relation between the hBN quality and the ODMR splitting in the limit of low irradiation fluence. Rather a key aspect is that ODMR splittings of irradiation created V–B close to 0 MHz were only reported in few-layer hBN [54]. Here, it was argued that not the purity of the hBN, but rather the 2D geometry plays a decisive role in reducing the number of nearby background charges resulting in a clean ODMR spectrum with almost negligible peak splitting.
@@ IV Conclusions
In summary, we quantified the defect creation within homogeneous hBN films using a helium ion microscope to site-selectively pattern of V–B defect ensembles. We analyzed the optical properties of the emitters as a function of ion fluence over more than two orders of magnitude. The spatially precise creation allowed us to study the magnetic response probed by ODMR contrast measurements in a consistent and systematic way. We simulated the level splitting obtained experimentally from the ODMR spectra employing a microscopic charge model, where the spatially random distribution of local charges results in a level splitting in ODMR. As a main finding, we observe a discrepancy between the scaling of the charge density, or equivalently the level splitting in ODMR, and the scaling of the defect PL, or equivalently the density of optically active defects, as function of ion fluence. Unlike other studies, which attributed the charge density in the microscopic model solely to V–B defects, we propose to include a background density in the model. With this modification, we were able to estimate the defect creation yield and compare it to molecular dynamics simulations. On one hand, further developments may involve the implementation of new theoretical models with the aim of addressing the fundamental question of the origin of the V–B charge [76], such as the nature of the charged state (dynamic or static) and the role of the local defect environment, as it has been done for N centers [77]. On the other hand, spatially precise creation of such emitters as demonstrated here may enable their integration in gated, light-emitting hetero-stacks, in which the PL emission can be tuned upon application of a bias voltage [78] and a tunneling current can be coupled to the in-gap defect states by means of graphene electrodes [60, 79].
@@ Acknowledgements.
We thank Moritz Fischer for helpful discussion and Dorte Rubæk Danielsen for help in the sample fabrication. N. S. thanks the Novo Nordisk Foundation NERD Programme (project QuDec NNF23OC0082957). Work at the Technical University of Munich was supported through the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Munich Center for Quantum Science and Technology No. (MCQST)-EXC-2111-390814868. C.K., L.G., J.F. and A.W.H. acknowledge funding through the International Graduate School of Science and Engineering (TUM-IGSSE, project BrightQuanDTUM). N.S. and M.W. acknowledge the support from the Independent Research Fund Denmark, Natural Sciences (project no. 0135-00403B) and from the Danish National Research Foundation through NanoPhoton - Center for Nanophotonics (project no. DNRF147). A.H. and I.B. acknowledge the NNF Biomag project (NNF21OC0066526) and the Danish National Research Foundation through the center for macroscopic quantum states (bigQ, project no. DNRF0142). A.V.K. acknowledges funding from the German Research Foundation (DFG), projects KR 4866/9-1, and the collaborative research center “Chemistry of Synthetic 2D Materials” CRC-1415-417590517. Generous CPU time grants from the Technical University of Dresden computing cluster (TAURUS) and Gauss Centre for Supercomputing e.V., Supercomputer HAWK at Höchstleistungsrechenzentrum Stuttgart are greatly appreciated.Conflict of Interest
@@
The data that support the findings of this article are openly available [80].
@@ A.2 Optical Characterization
Optical properties are characterized by diffraction-limited confocal scanning photoluminescence (PL) mapping, Raman microscopy, and time-resolved photoluminescence (TRPL) under ambient conditions. PL measurements are performed on a custom confocal microscope featuring a 532-nm laser (Cobolt, Hübner GmbH). The objective is a Nikon TU Plan Fluor /0.9 NA EPI D. Typical laser powers used for the measurements are around 3 mW (0.122 mW measured at the entrance aperture of the objective). The signal is collected by a spectrometer (Andor-Solis SR-303i, 150 grooves/mm grating with blaze at 800 nm) equipped with a CCD (Andor Newton). Before the He-ion beam irradiation described below, we verify that the selected hBN flakes contain low background luminescence in the spectral range where emission from V–B is expected ( – ) as well as that only a few intrinsic quantum emitters, with typical emission lines in the visible, are present. Raman spectroscopy is carried out in a commercial Raman microscope (WiTec Alpha, 532 nm laser, cw-excitation, 300 grooves/mm grating with blaze at 500 nm, Zeiss EC Epiplan-Neofluar objective lens / 0.9 NA).
@@
Time-resolved PL measurements (Figure A.3) are performed on the same setup as for PL experiments. For the excitation, we use a PicoQuant 515 nm laser diode in pulsed mode (200 µW of power, 10 MHz repetition rate), in combination with an avalanche photodiode (MPD) and a Time Tagger 20 (Swabian Instruments). The width of the instrument response function (measured on a reflective substrate) is around 250 ps. For fitting, we use the Python package Lifefit [81], which uses an FFT-based, iterative reconvolution of exponential decays with an instrument response function to model the lifetime data. During fitting, the amplitude weights of the individual decays are optimized using a non-negative least squares algorithm, while the lifetime parameters are optimized by non-linear least squares regression. To account for shot noise, Lifefit uses as weights in the least-square optimization, where is the number of photon counts in a time bin.
@@ A.3 Helium-ion irradiation
To create luminescent defects in the hBN films, we use a focused He-ion beam in a helium-ion microscope (Zeiss Orion). As irradiation patterns, we choose arrays of circular patches with diameter and 2 µm pitch. The former is on the order of the typical spot size in the confocal microscope, such that He-irradiated spots appear point-like in the PL images (Figure A.2), and the latter is much larger, so that the exposure spot can be well separated laterally. We use an acceleration voltage of , a beam spacing of 5 nm in x- and y-directions, a beam aperture of , dwell time of and a beam current of . The values of the irradiation fluence are set in on the HIM (see Table 1 in the main text for the fluences employed). To minimize unwanted irradiations, we use a large field of view and low imaging fluence for locating the heterostructures. For hBN placed directly on Si/SiO2, we find significant background PL after ion exposure. The PL originates from both the Si/SiO2 substrate and the hBN (Figure A.4). To avoid such background PL, all the measurements in the main manuscript are conducted on hBN/graphite/SiO2/Si heterostructures, resulting in spectra virtually free from unwanted background luminescence.
@@ A.6 Molecular Dynamics (MD) simulations
MD simulations are carried out using the LAMMPS package [62] to extract the statistics of defect creation produced by the interaction of He ions with hBN. The interaction of the ion beam with the hBN structure is modeled via an extended Tersoff potential for boron nitride (BN-ExTeP) [83], with a smooth transition to the Ziegler–Biersack–Littmark (ZBL) potential [84] for small distances. The simulation spans a material thickness of and is performed for a reference fluence, given by a number of 990 ions and a covered surface of 35 Å 35 Å. The MD approach is compared to other Monte Carlo methods, which are the TRIDYN code [85] and the SRIM software [86].
A.7 Comparison of ODMR splitting across different studies
In Table 3 we compare our ODMR results with previously reported studies that explicitly observed V–B defects and provided sufficient information about the hBN quality, the irradiation parameters and the ODMR results.
| Reference | Splitting | hBN source | Irradiation | Fluence |
| [MHz] | (thickness [nm]) | Method | [× 1015 cm-2] | |
| This work | HQgraphene (80) | FIB He+ (30 keV) | 1 – 30 | |
| Gottscholl, Nat. Mater. 2020 [6] | HQgraphene (80000) | Therm. neutrons | 2300 | |
| Guo, ACS Omega 2022 [21] | HQgraphene (10 - 100) | IB N+ (30 keV) | 0.01 - 1 | |
| HQgraphene (10 - 100) | IB He+ (30 keV) | 0.1 | ||
| Ren, J. Lumin. 2023 [23] | HQgraphene (>1000) | FIB He+ (50 keV) | 0.1 | |
| Gong, Nat. Commun. 2023 [44] | Commerc. avail. hBN | IB He+ (3 keV) | 0.03 - 1 | |
| Gao, Nano Lett. 2021 [19] | 2D Semiconductors | IB He+ (0.3 - 2.5 keV) | 0.05 | |
| Zabelotsky, Appl. Nan. Mat. 2023 [51] | 140 | NIMS (100/150) | FIB N+ (12 keV) | 0.625 |
| Sasaki, Appl. Phys. Lett. 2023 [50] | NIMS (66) | FIB He+ (30 keV) | 1 | |
| Suzuki, Appl. Phys. Expr. 2023 [52] | 150 | NIMS (>100) | FIB N+ (40 keV) | 1 |
| 90 (high temp. irr.) | NIMS (>100) | FIB N+ (40 keV) | 1 | |
| Liang, Adv. Opt. Mat. 2023 [28] | NIMS (>100) | FIB He+ (500 keV) | 1 – 100 | |
| Udvarhelyi, npj Comput. Mater. 2023 [55] | Own prod. (bulk) | Therm. neutrons | 26 - 260 | |
| Durand, Phys. Rev. Lett. 2023 [54] | Own prod. (15) | Therm. neutrons | 26 | |
| Own prod. (few layer) | Therm. neutrons | 26 | ||
| Haykal, Nat. Commun. 2022 [56] | 120 | Own prod. (bulk) | Therm. neutrons | 26 - 260 |